### v1

# Evaluating LLMs on microservice-based applications: how complex is your specification?

Abstract
In this paper we evaluate how far LLMs have advanced in generating code for real-world problems. Specifically, we explore code synthesis for microservice-based applications, a widely used architecture pattern. We define a standard template for specifying these applications, and we propose a metric for judging the difficulty level of a specification. The higher the score, the more difficult it is to generate code for the specification. We develop a framework to automate the process of testing the LLM-synthesized code for a microservice using unit tests. Our experimental results show that strong LLMs (like GPT-3o-mini) do fairly well on medium difficulty specifications but do very poorly on those of higher difficulty levels. This is due to more intricate business logic, a greater use of external services, database integration and inclusion of non-functional capabilities such as authentication. We analyzed the errors in LLM-synthesized code and report on the key challenges LLMs face in generating code for these specifications thereby suggesting future research directions to improve code synthesis for real-world problems.
1 Introduction
The rapid maturation of Large Language Models (LLMs) for code generation has widely influenced software development, with LLMs now routinely used for software engineering tasks including code generation. Benchmarks have become the de facto standard for measuring an LLMs’ performance. Early benchmarks, such as HumanEval[2] and MBPP[1] determine how well LLMs perform in generating a single function of increasing algorithmic complexity. State-of-the-art models (including reasoning models such as GPT-o1 and GPT-o3 and DeepSeek-R1) show impressive results against these benchmarks, with results approaching 100% for HumanEval [8].
However, these benchmarks shed little light on how well these models can perform on real-world problems, where the complexity is of a different nature. In real software development, complexity expresses itself not only in the computation (algorithm) of a particular component, but in the integration of different software components and in the distributed nature of modern systems; i.e., Cloud Native architectures [21]. Newer benchmarks, such as SWE-Bench [10] have been developed to address these more realistic scenarios by asking LLMs to fix a specific bugs in real GIT repos [33, 17, 31, 3] or to synthesize code to add a new feature/function to a repo [13, 32, 15].
The main challenge that GIT repo benchmarks poise to LLM code generation is the size of the context. Popular GIT repos often have hundreds of thousands LOC [10] which is too large to fit into LLM context windows. Hence newer techniques, including agentic solutions, often include information retrieval techniques and focus on how to extract the right context to feed the LLM [30, 28, 29].
There is an orthogonal problem in generating software for real-world problems that has received less attention; namely, how to specify a complete solution to be generated. Real-world software is complex and requires far more descriptive text than required for fixing a GIT repo issue. How well do LLMs perform in this scenario?
To make the problem concrete, we focus on LLM code synthesis for a core design pattern of modern systems, Microservice-Based Applications (MSBAs). These apps have the following characteristics:
  • Microservices (MSs) as the architectural pattern for system decomposition [20], usually deployed in containers such as Docker.
  • RESTful APIs as the core communication mechanism [4].
  • Use of 3rd party packages and external services to easily incorporate existing functionality.
The GitHub repo problem focuses on code generation where the code base is complex and cannot fit into the context window of the LLM. The MSBA problem, in contrast, focusses on code generation from scratch, where the specification can easily fit into the context window of the LLM, but the challenge is in the complexity of the spec; i.e., the intricacies of the software being developed, not code size. For instance, SWE-Bench issues provided to an LLM to fix are 195 words on average [^1] However, the max length for SWE-Bench issues is 4477 words. [10]. Our MSBA examples have descriptions of 1399 and 1905 words. While these are not full-fledged production-level apps, they serve as a good testbed for our experiments. In Section 3.2 we discuss metrics for measuring the complexity of MSBAs. As far as we know, there are no existing benchmarks for MSBAs, and little if any work has examined using LLMs to generate such systems.
The purpose of this paper is to provide a preliminary study on how well LLMs perform in this important domain, and to encourage more research to address the challenges we find.
2 Challenges in generating MSBAs
Some of challenges in synthesizing MSBAs from LLMs are:
Integration: includes reconciling type and naming mismatches between different frameworks and libraries.
History and exceptions: the response to a REST request can depend on the history of previous requests, and depending on that history, a request could generate a normal result status code and return value, or one of many exceptional (HTTP) status codes.
Non-functional considerations: Orthogonal non-functional considerations, such as authentication or logging, complicate code generation.
Dependencies: A single MS can depend upon many other MSes and 3rd party services [13]. This also requires understanding the right packages and versions to use.
Distribution: MSBAs are inherently distributed systems and must deal with complexities inherent in such systems.
Specification: There is no standard for specifying a MSBA to an LLM. Unlike simple functions, a docstring and a signature do not suffice.
On the other hand, MSBAs using REST services often follow a well-established template (see Appendix C for a small example) giving hope that LLMs can exploit this template to generate correct code.
The research questions we seek to answer are:
  • How should one specify a MSBA to facilitate LLM code generation and how can one measure the complexity of a specification? This is discussed in Section 3.
  • How to test an MSBA? We address this in Section 4.
  • How well do LLMs perform in generating MSBAs? How does the complexity of the specification affect LLM performance? Our experimental results are given in Section 5.
3 Specifying a MSBA
To describe an MS that is part of a MSBA you need to identify:
  • The resources that are part of the interface. These are the “nouns” of an MSBA, the things that get operated upon. For instance, the resource /books is a collection of books, while /books/{id} is a specific book.
  • The requests (HTTP verbs) that can be issued on these resources. The basic verbs are the CRUD-like requests POST, GET, PUT and DELETE. Each operation will specify additional details. For instance, a GET request must specify its payload (parameters), allowable query strings, and details specific to HTTP and RESTful APIs, such as media types accepted and status codes to be returned. The spec may also limit whom is allowed to issue requests.
  • The computations that must be performed and invariants that need to be maintained when handling a request.
  • Dependencies such as external services required to process requests and how to access their API keys.
{"Name": "Cardholders microservice","Description": "Responsible for creating, retrieving, updating, and deleting     cardholder accounts. This service also provides the ability to retrieve     how much money in fines is owed by a specific cardholder.”"Resources": "/cardholders, /cardholders/{id},/cardholders/fines/{id}. An id is     of type string. Each cardholder (/cardholders/{id}) consists of a JSON     document containing the field names ‘name’, ‘email’, and ‘id’. The values     of each of these is a JSON string. The resource /cardholders/fines/{id}     consists of a JSON document with a field ‘fineAmount’, whose value is     of type float, and a field ‘id’, of type string.","REST requests": "GET and POST request on /cardholders. GET returns a     JSON array of all records in the resource. The GET request may specify a     query string of the form ‘field=value’. In this case, only those cardholder     records satisfying that constraint should be returned. The POST request     provides a JSON document containing the name and email and returns the     id for the newly created record.     GET,PUT and DELETE requests on /cardholders/{id}. The GET request     returns a JSON record of the requested resource.  The PUT request provides     a JSON payload of the updated JSON resource’s fields and returns the id of     the updated resource. The DELETE request returns the status of the request     with no content.     GET request on cardholders/fines/{id}.  The request returns the JSON     document representing this resource.","Additional details”: "The fine amount owed is computed as follows: at any given     day, the fine per overdue book is equal to number-of-days-overdue *     FINE_PER_DAY, where FINE_PER_DAY equals \$.50. The cardholder owes     the sum of all fines for all books he borrowed. To compute this amount,     this service needs to invoke the Borrows MS to retrieve information on books     borrowed by this cardholder.","Deployment": "The Cardholders MS runs in a Docker container and listens on     external port 5001. Another MS can send it requests using the URL     http://cardholders:5001/cardholders."}
Figure 1: Cardholders microservice
In Figure 1 we illustrate the specification of a Cardholders MS, part of a simple public library application we present in Section 5.3. “Description” provides the purpose of the MS. “Resources” lists the REST resources that the service is responsible for. “REST requests” defines the type of REST requests supported for each resource, and how resources are to be represented. “Additional details” provides functional requirements for the service. “Deployment” states deployment details; for instance, that the micro-service will run in a Docker container, the port the MS listens on and the URL used to send this MS requests. An alternative to this format would be to use OpenAPI (Swagger) specifications [18] but we would need to add information necessary to implement the server implementation. The format we provide is more compact.
All prompts we use are given in Appendix A, including the LLM Code Generation Guidelines. These guidelines instruct the LLM to build the MSes using REST services that execute Python code, and that store data in a MongoDB. It also defines which HTTP status codes should be returned in case of normal or exceptional execution.
3.1 Creating MSBA specifications
To evaluate LLMs on MSBAs, we created 8 MSes across 2 MSBAs. We created our own MSBAS for two reasons. First we want to make sure that the LLMs have not been contaminated - have not seen these MSes before. Second, we want to compare LLM’s code generation capabilities on MSBAs of increasing specification complexity, but still small enough to focus on core MSBA challenges (Section 2) and not on extraneous issues. We built the MSBA specifications using the following process:
  1. 1.
    We wrote the spec of each microservice as part of a complete MSBA specification, and reviewed it with an LLM for completeness and consistency.
  2. 2.
    We wrote a Ground Truth (GT) implementation for the spec.
  3. 3.
    We wrote unit test cases based upon the spec and used them to test the GT solution.
  4. 4.
    When testing and finding an error, we checked the cause of the error. (a) If it was a mistake in the GT implementation, we fixed the GT implementation. (b) If the test case was not faithful to the spec, we corrected the test case. (c) If the spec was ambiguous or incorrect, we fixed the spec.
  5. 5.
    We repeated step 4 until the GT implementation passed all tests.
  6. 6.
    We used the spec to generate preliminary code from an LLM and tested that code with the unit tests. We discovered a few more errors this way, and repeated until we were confident in the correctness of the GT implementation and unit test cases.
The specifications and list of unit tests for each MSBA can be found online222 https://github.com/danny-m-yellin/LLMs4MSBA-artifacts.
3.2 The complexity of an MSBA specification
We measure the complexity of a a MSBA in several ways:
  1. 1.
    Size: the number of words in the MSBA specification. Although we generate each MS one at a time, we provide the entire MSBA specification to the LLM because it often needs the entire spec to generate even a single MS correctly.
  2. 2.
    Dependencies: Let $M$ be a MS and $E$ be a service endpoint. $M$ is dependent on $E$ if $M$ invokes $E$. $E$ may be another MS in the application or it may be an external service. We count the number of dependencies for each MS in the MSBA, and the sum of these is the dependency complexity of the MSBA. We use the GT code to determine the number of dependencies. Note that if $M$ invokes $E$ from multiple places in the code, it is still counted as a single dependency.
  3. 3.
    Packages is the number of packages the MS requires. We count the number of packages that the GT implementation imports.
  4. 4.
    LLM-as-a-Judge is the score, between 1 and 5 assigned by the LLM to the difficulty of implementing the spec.
The specification size indicates the number of details the LLM must contend with. The number of dependencies and packages indicates the amount of external knowledge required by the LLM. These metrics are not exact; its possible to have a large specification that is simple to implement and a small specification that is difficult to implement. Similarly not all external dependencies are equal; some express more intricate interfaces than others. Nonetheless, these metrics give a rough estimate on the complexity of an MSBA spec and are easy to compute. The LLM-as-a-Judge serves as a validation for the other metrics. Furthermore, the LLM explains why it gives the score, which provides further insight into the complexity of the spec.
4 Testing MSBAs
The result of a request to a MS is sensitive to history. For instance, the outcome of a GET request is dependent on the history of the preceding POST, PUT and DELETE requests. One needs to test multiple scenarios for a GET to determine whether it functions correctly. Furthermore, the unit tests must also check that the appropriate exceptions are raised when warranted; e.g., when a GET request is issued on an item previously deleted, the status code 404 must be returned whereas if the request provides the wrong media type a status code 415 is returned. Developing adequate unit tests can be time consuming and tedious. To help simplify and make testing of REST APIs more consistent, we developed a REST API testing framework, that consists of 4 modules.
The main module is (i) Unit Test, which issues the unit tests. It relies on the (ii) HTTP Service Controller to format and invoke the requests according to the HTTP protocol. We divide unit tests into two sets: common unit tests and MS-specific unit tests. Common tests are those where the result depends on the pattern of the requests, without concern for their specific input and output values. For instance, if we issue requests POST $X_{1}$, POST $X_{2}$, POST $X_{3}$, DELETE $X_{2}$ and finally GET, we expect the return value to be the JSON array containing $X_{1}$ and $X_{3}$ and nothing else. This is true independent of the actual values of $X_{i}$. We rely on this insight to parameterize unit tests. Parameter values, expected returned values, and any other MS specific values needed to test the MS are defined in the (iii) Database module. The Common Unit Test module fetches the appropriate values from the DB when testing a MS; for the example above, each MS will provide different values for $X_{1}$, $X_{2}$ and $X_{3}$.
The common unit tests may still require some customization, and not every custom test applies to all MSes. However, the framework greatly simplifies building tests for a new MS. Likewise, we created a reusable parameterized (iv) assertion library that checks return values satisfy assertions. For example, the assertion assert-fields-contains-value(response: requests.Response,f:str,v) asserts that the response is a JSON array, each element in the array is a struct with field f, the field f is itself a JSON array, and f contains the value v.
In addition to common unit tests, we need to test MS-specific functionality. For example, we want to check that the Cardholders MS (Figure 1) correctly computes the fine for overdue books. MS-specific tests need to be customized per MS.
Say that an MSBA is made up of 4 MSes: $M_{1}$, $M_{2}$, $M_{3}$, and $M_{4}$. Let $M^{GT}_{i}$ be the GT implementation of $M_{i}$ and $M^{LLM}_{i}$ be the implementation of $M_{i}$ generated by the LLM. To test $M^{LLM}_{2}$ we run the MSBA using the MS implementations $M^{GT}_{1}$, $M^{LLM}_{2}$, $M^{GT}_{3}$, and $M^{GT}_{4}$ and then test the application. In this way we isolate the test failures due only to $M^{LLM}_{2}$.
5 Experiments
5.1 Methodology
Given an MSBA specification, we ask the LLM to generate the MSs of that MSBA. Our implementation is a workflow (WF) built on top of LangGraph [11], as illustrated in Figure 2. After (1) initializing the WF, it (2) prompts the LLM with the MSBA description including code generation guidelines, asking it to generate one MS $M$. The prompt is given in Appendix A. It then (3) creates a requirements.txt file for the generated code. (4) Next it sets up the Docker containers to execute the LLM generated code $M$ and the GT implementations for all other MSes and then (5) executes a Docker Compose script to execute the unit tests.
Figure 2: Workflow for generating and testing an MSBA. The blue arrows indicate control flow. The dotted black arrows indicate writing and reading data. Blue rectangular boxes are LLM interactions.
(6) Once the tests complete, or the execution is interrupted by run-time errors, the containers are shutdown and (7) the error messages are examined. If there are no runtime and test errors, the LLM synthesized code is successful and the workflow (11) records the results. (8) Otherwise it checks if it has reached the maximum number of code-generation trials. If not, (9) the WF asks the LLM to reflect [23] on the errors and produce a plan to fix the generated code to prevent the top 5 errors. This prompt is given in Appendix A. Because of context window limitations of LLMs, and because the runtime may produce a very long trace of each error, we use a script to remove parts of the error messages that are not useful to debug the issue. For some LLMs, we need to further limit the size of the error messages to fit into the context window. (10) The original code together with the result of the reflection is then provided to the LLM, which is instructed to generate code that fixes the problems. This prompt is given in Appendix /refappendix-prompts. The workflow then loops back to step (3). The entire WF repeats for each MS.
In summary, our workflow automates the entire process from generation to testing to feedback by invoking multiple agents. It uses LangGraph to orchestrate the flow, runs the generated code safely in Docker containers, and uses reflection to improve results.
5.2 Installing packages
We run the LLM generated code in a Docker container. The Dockerfile for that container contains a command to install dependencies in the requirements.txt file using pip. We generate the requirements.txt file by scanning the generated code for import statements [^3] In our prompt we asked the LLM to document which packages to install. But we found that it would make mistakes and it was a better solution to install packages based upon the import statements. . This was problematic as pip would fail when installing a package that comes with the standard python distribution (such as uuid or json). To get around this problem, we omitted packages included in the Python standard library.
In the restaurant application (Section 5.4), the generated code for the authentication MS would always import the jwt package, but it required the installation of pyJWT. (To its credit, the LLM would note this in a comment.) For this and other similar issues, our code substitutes the package name in the import statement with the correct package. Nonetheless, as we show below and documented by other research [24, 12, 9], choosing the right package and version was still problematic.
5.3 The Public Library Application
Our first MSBA is a public library application comprised of 4 MSs: Cardholders, Books, Borrows, and Logs. The Cardholders service was given in Figure 1. Books is responsible for managing the books in the library, Borrows is responsible for tracking books on loan, and Logs keeps a record of each book that is borrowed from the library. The specification of these MSes can be found on-line (see footnote 2).
We test this application using 3 different LLMs: GPT-3.5-16K, GPT-4o-mini, and GPT-o3-mini. For each LLM we performed two sets of experiments, one using a 0-shot prompt, and one using a $1$-shot prompt [5]. The $1$-shot prompt, given in Appendix C, provides example code using Flask and MongoDB for a small MS. For each LLM and each MS of the MSBA we repeat each experiment 15 times. For GPT-3.5 and GPT-4o-mini we perform 5 experiments for each temperature $0.0$, $0.3$, and $0.5$. GPT-o3-mini did not support different temperature settings when we ran our experiments and hence all experiments used the default temperature. Each experiment tests the MSBA with unit tests for each MS. There are 15 unit tests for Cardholders, 14 for Books and Borrows, and 7 for Logs. The latter has fewer tests because Logs only allows external actors to perform GET requests; they cannot create, delete or modify logs on their own. For each experiment, if the LLM generated code does not pass all the unit tests, we prompt the LLM to reflect on the nature of the errors and then ask it to regenerate code, as explained above. We only do this once (MaxGen = 2 in Figure 2). We refer to the initial generated code as V0 and the regenerated code as V1.
5.3.1 Results
Tables 1 and 2 shows the results of our experiments on GPT-3.5 with the 0-shot and $1$-shot prompts respectively. The $1^{st}$ column is the name of the MS. As not all code generated by the LLM was executable, the $2^{nd}$ column lists the fraction of the 15 experiments that generated executable V0 code permitting execution of the unit tests. The $3^{rd}$ column lists the number of experiments that generated executable V1 code. The $4^{th}$ and $5^{th}$ columns present the percent of unit tests that passed using V0 and V1 code respectively, averaged over all experiments that produced executable code444 We do not consider unexecutable generated code in this percentage. We wanted our results to focus on how often executable code was functionally correct. . The $6^{th}$ column is of the form $n_{1}$-$n_{2}/t$, where $n_{1}$ ($n_{2}$) is the highest number of unit tests passed in any experiment for V0 (V1) and $t$ is the total number of unit tests for that MS. The last column is the percent of experiments in which the number of unit tests passed by V1 was greater than the number of unit tests passed by V0, considering only those experiments in which V1 was executed.
Name # V0 # V1 # V0 % V1 Highest % V1
testable testable passed passed # passed improved
Cardholders 13/15 13 26.4 50.2 # 7-8/15 76.9
Books 15/15 15 38.1 37.1 # 7-7/14 13.3
Borrows 15/15 15 28.1 32.4 # 7-7/14 33.3
Logs 15/15 15 30.5 26.7 # 4-4/7 26.7
Table 1: Library MSBA results using GPT-3.5 with 0-shot prompt
Name # V0 # V1 # V0 % V1 Highest % V1
testable testable passed passed # passed improved
Cardholders 13/15 13 59.3 78.4 # 13-13/15 61.5
Books 15/15 15 61.9 66.2 # 13-13/14 40.0
Borrows 15/15 15 87.2 72.4 # 13-14/14 6.7
Logs 15/15 8 83.8 62.5 # 7-6/7 12.5
Table 2: Library MSBA results using GPT-3.5 with 1-shot prompt
Tables 3 and 4 present the results for GPT-o3-mini using the 0-shot and 1-shot prompts.
Name # V0 # V1 # V0 % V1 Highest % V1
testable testable passed passed # passed improved
Cardholders 13/15 5 94.4 0.0 # 15-0/15 0.0
Books 13/15 12 83.5 16.7 # 14-14/14 16.7
Borrows 15/15 10 85.2 20.0 # 14-14/14 20.0
Logs 14/15 4 90.8 17.9 # 7-5/7 25.0
Table 3: Library MSBA results using GPT-o3-mini with 0-shot prompt
If V0 code passes all unit tests, V1 code would not be generated. Hence sometimes there were very few V1 samples generated. If these samples have a major flaw (like uncommented text in the code that caused the entire MS to produce runtime errors), it resulted in the $0\%$ passing value in the $5^{th}$ column.
Name # V0 # V1 # V0 % V1 Highest % V1
testable testable passed passed # passed improved
Cardholders 13/15 4 89.2 0.0 # 15-0/15 0.0
Books 15/15 14 81.4 21.4 # 14-14/14 21.4
Borrows 15/15 13 87.6 30.8 # 14-14/14 26.7
Logs 15/15 5 90.5 0.0 # 7-0/7 0.0
Table 4: Library MSBA results using GPT-o3-mini with 1-shot prompt
We omit the results for GPT-4o-mini due to space considerations. In general it performs better than GPT-3.5 but not as good as GPT-o3-mini. See Appendix D.
5.3.2 Discussion
Looking at column 3 of Table 1 we see that GPT-3.5 performs poorly with the 0-shot prompt, never reaching 40% average test success. In contrast, in Table 3 we see that GPT-o3-mini performs quite strongly, averaging between 83.5 and 94.4%. Keep in mind that this is the average number of unit test cases passed; it does not mean that 94.4% of the instances pass all test cases. While GPT-3.5’s performance increases upon reflection (V1 code), it still lags significantly behind GPT-o3-mini. Furthermore, from column 6 of Table 3 we see that some GPT-3o-mini V0 samples for each service pass all tests. This affirms the strength of GPT-o3-mini in generating code.
A second observation, based upon columns 5 and 7 of these tables, is that reflection is effective for GPT-3.5 but much less so for GPT-o3-mini. We attribute this to two factors. First, GPT-o3-mini was often successful in passing all unit tests using V0, and hence there were fewer V1 experiments. Second, even when V0 did not pass all tests, it had a high success rate, so it created a higher bar for V1 to improve upon.
A third observation, based upon column 4 of Tables 2 and 4, is that providing an example (1-shot prompt) significantly improves performance of GPT-3.5 but mostly degrades performance with GPT-o3-mini. This is often the case with reasoning models like GPT-o3-mini [^5] For instance, the creators of DeepSeek -R1 report that “When evaluating DeepSeek-R1, we observe that it is sensitive to prompts. Few-shot prompting consistently degrades its performance. Therefore, we recommend users directly describe the problem and specify the output format using a zero-shot setting for optimal results.” [7] See also [27]. , and especially in our case, where performance is high without the example, and the example can cause overfitting.
Looking at the generated code, we find that GPT-3.5 often makes mistakes in the basic REST patterns and requires the example to nudge it to implement the REST APIs as specified. In contrast, GPT-o3-mini understands this pattern well and does not require much assistance.
A general comment is warranted on these statistics; namely, not all unit tests are equal. Some unit tests may require complicated logic while others just check very simple REST patterns that do not even require knowledge of the MS. For instance, one of the unit tests checks that a POST request with a non-JSON payload fails. Hence while percentage of unit tests passed still gives us an overall idea on the strength of the LLM, we require a deeper analysis examining which tests pass or fail to obtain more insight (see Section 5.7).
With respect to temperature, while the $0.3$ setting often performs best for V0 and $0.5$ for V1, the differences in performance between different temperatures is usually less than $10\%$. See Appendix D for details.
5.4 The Restaurant Application
The 2nd application is a restaurant nutrition MSBA comprised of 4 MSs, Authentication, Dishes, Profile, and Ratings. Authentication is responsible for authenticating users, including registration, login, and token validation. Dishes is responsible for creating, retrieving, updating, and deleting dishes offered by the restaurant. Only the restaurant owner (a predefined role) is allowed privileged operations such as POST, while general users are only allowed to perform GET requests. This service invokes the 3rd party API Ninjas nutritional service, available at https://api-ninjas.com/api/nutrition, to provide nutritional information on the dish. Profiles is responsible for creating, retrieving, updating, and deleting user profiles. Each user profile contains a nutritional diet for that user, and allows the user to retrieve only those dishes that satisfy his/her diet. Ratings is responsible for maintaining ratings of restaurant dishes and computing the average rating for a dish. All users can rate a dish but a user is only allowed to rate a dish once.
Given the poor results of GPT-3.5 on the Library app, we use only GPT-4o-mini and GPT-o3-mini for our experiments on this MSBA. For similar reasons, we only experiment with a 0-shot prompt. For GPT-4o-mini we performed 20 experiments for each MS. For the ratings MS, 8 of the experiments were at temperature $0.0$, 7 were at temperature $0.3$ and 5 were at temperature $0.5$. For all other microservices, there were 10 experiments at temperature $0.0$, and 5 experiments at each of the temperatures $0.3$ and $0.5$. For GPT-o3-mini we performed 15 experiments for each MS. As mentioned above GPT-o3-mini did not support different temperatures at the time of our experiments, so we used the default temperature.
Each experiment consists of 9, 15, 10 and 8 unit tests for the Authentication, Dishes, Profile, and Ratings MSes respectively. As above, for each experiment, if the LLM generated code does not pass all unit tests, we ask the LLM to reflect on the nature of the errors and then regenerate code.
5.5 Results
Tables 5 and 6 give the results of testing on GPT-4o-mini and GPT-o3-mini LLMs. We see that their performance degrades significantly on the restaurant MSBA compared to the library MSBA.
Name # V0 # V1 # V0 % V1 Highest % V1
testable testable passed passed # passed improved
Authentication 0/20 20 0 57.7 # 0-6/9
Dishes 20/20 20 15.3 23.7 # 5-7/15 55.0
Profile 20/20 19 3.5 8.4 # 1-1/10 47.3
Ratings 20/20 20 1.3 0.6 # 1-1/8 5.0
Table 5: Restaurant MSBA results using GPT-4o-mini with 0-shot prompt
Name # V0 # V1 # V0 % V1 Highest % V1
testable testable passed passed # passed improved
Authentication 13/15 11 65.8 79.3 # 8-9/9 81.8
Dishes 15/15 15 23.6 25.8 # 5-6/15 20.0
Profile 15/15 15 14.0 15.3 # 2-2/10 26.7
Ratings 15/15 15 10.8 39.2 # 3-6/8 66.7
Table 6: Restaurant MSBA results using GPT-o3-mini with 0-shot prompt
5.6 Description Complexity
Before analyzing the results of these experiments, we first examine the complexity of the restaurant and library MSBAs, given in Table 7. The description size is for the 0-shot prompt. The 1-shot prompt adds $490$ words.
The Restaurant MSBA has a description size about .36.36 larger than the Library application, and has twice the number of dependencies, although the package complexity is the same. This at least partially explains the difference in LLM performance between the library and restaurant MSBAs.
Name Description # Dependencies Avg # Packages LLM assigned
size (words) per MS difficulty
Library 1399 3 4 3
Restaurant 1905 6 4 4
Table 7: Description complexity of MSBAs
The LLM-as-a-Judge score, provided by Claude Sonnet 4, confirms this conclusion. It assigns the library application a difficulty of 3 (out of 5) while the restaurant application received a complexity score of 4. It says that the library app has “moderate complexity” and that it is not overly complex because of the “Clear, well-defined requirements”, “Standard CRUD operations”, “Common technology stack”, “Simple data models”, and “Limited business logic”. On the other hand, factors that add complexity include “Inter-service communication”,“External API integration”, “Business rule enforcement”, “Date calculations”, “Error handling”, and “Distributed data consistency”.
For the restaurant MSBA, the LLM states the following “High Complexity Elements” contributing to the 4 rating: “Multi-service Architecture”, “Authentication & Authorization”, “External API Integration”, “Complex Business Logic”, “Database Operations”, “Error Handling”, and “Query Processing”. It did not give it a 5 rating because of “Well-defined specifications with clear requirements”, “Standard technologies (Python, MongoDB, REST APIs)”, “No real-time processing or complex distributed systems patterns” and “No advanced security beyond basic token auth”. In Appendix B we provide the LLM-as-a-Judge prompt and the complete comments provided by the LLM to explain its score.
5.7 Discussion
From Table 5 we see that GPT-4o-mini does extremely poorly on the restaurant application, and while GPT-o3-mini (Table 6) does better, it still never performs above $25\%$ passing unit tests with V0 code except for the Authentication MS, where it achieves $65.8\%$. We also note that neither V0 nor V1 code ever achieves a perfect score, with the exception of the Authentication service. The V1 code of that service, as seen in column $6$, sometimes passes all $9$ unit tests. This is true for 6 out of the 11 V1 testable implementations.
To get a deeper understanding of the cause of the errors, we examined the run-time error-logs, the GPT-o3-mini LLM-generated reflection results, as well as the generated code. Some of the persistent problems we found in the code are the following:
  • Incorrect package usage. For instance, even though we installed pyJWT (see Section 5.2 ), the authentication code would sometimes use this package incorrectly. This might be due to an incorrect version being installed. Along similar lines, Ratings would consistently import the flask_jwt_extended package. A typical error message was the following:
    **Error**:‘ImportError: cannot import name ‘jwt_optional’ from ‘flask_jwt_extended’
    We tried multiple versions of the flask_jwt_extended package on the code to see if this would fix the problem, but none corrected the problem, or would result in other runtime errors. In short, incorrect package usage is often a fatal problem in the generated code [24, 12].
  • Superficial understanding of APIs. This is best illustrated by the Dishes MS using https://api-ninjas.com/api/nutrition. The spec purposely changes the names of the nutritional elements used in the MS from those provided by the API. Furthermore, for some nutritional elements, such as protein, the API returns the string “Only available for premium subscribers”. These elements are not required by the app, but the code must know that not every element in the array will be numeric or it throws an exception. This level of API understanding was beyond the LLM. To overcome this problem, we developed a new prompt that explicitly states the details of the API. Although using this prompt the generated code was able to address these API complexities, the test results remained largely the same due to other issues, like those enumerated below. Finding appropriate APIs and correctly implementing them is another recognized LLM issue [14, 16, 22, 25].
  • Lacking or incorrect datatype conversions. For instance, the Dishes implementation checks that the type of a parameter is a list. But it first needs to use json.loads to convert the string data to JSON. Its failure to do so results in a runtime error and the POST request fails. This has a cascading effect - since the POST fails, subsequent GET requests fail. This error is pervasive and contributes to the very low passing rate for some of the services.
    A similar issue is converting between REST APIs parameters, passed as a string, and the type required to use in an API invocation. For example, a resource (e.g., a specific dish) is identified by its string ID in the REST API. But MongoDB uses, by default, ObjectIDs to identify Mongo documents. Failure to convert between these datatypes was a common error in the library application using GPT-3.5. GPT-o3-mini largely understands the need for this conversion and avoids these mistakes. However there are still times when it fails to perform the conversion due to subtleties of the Mongo API. Specifically, the generated code: dishes_collection.insert_one(dish_record) followed by jsonify(dish_record) causes a run-time error:
    TypeError: Object of type ObjectId is not JSON serializable. This is because by default Mongo adds an ObjectId to the dish_record. One needs to remove this added field or convert it to a string before applying the jsonify() method. Mistaken LLM type inference is a problem for LLMs [19].
  • Disregarding specification details. A good example of this problem is wrong return codes. In the prompt we state “If an authentication token is not provided or is invalid, then 401 (Unauthorized) is returned.” However, the generated code returns status 400 (Bad request). This same sort of error occurs for other return codes as well. Providing good error messages to the LLM on why a particular unit test fails will allow it to correct the problem. In this case, the LLM V1 implementation was able to fix faulty return codes.
6 Related Work
Extending LLM code generation to tackle more realistic real-world problems for software engineers is an active area of research. One trend in this direction is software engineering agents, who facilitate developers in their everyday tasks. More specifically, as related to our work, is research into extending LLM code synthesis to tackle more realistic and complex scenarios. Much of this research is focussed on automatically correcting issues or extending the functionality of a solution in a GitHub repository [10, 13, 3, 30, 15]. While the fix or code extension may be relatively minor, the difficulty is in providing the LLM with the correct context, which can be non-trivial given the size of the repository. In contrast, our work focusses on taking a specification of an application that has multiple microservices, and generating code from scratch to realize the specification. The challenge in this problem is the complexity of the specification. It requires generating code that incorporates business logic, invokes the appropriate internal and external dependencies, incorporates non-functional considerations such as logging and authentication, and handles a variety of inputs and multiple error conditions.
We found that for moderately difficult specifications GPT-o3-mini did quite well. However, for more difficult specifications it performed quite poorly. By examining the errors, we found specific issues that it had trouble resolving. For instance, one problem is that LLMs will sometimes hallucinate when importing packages [24, 12] . This displays itself when the LLM imports a package that does not really exist, or invokes non-existent APIs of in a package. Locating dependencies and implementing the APIs correctly is another recognized problem [14, 16, 22, 25].
While other research has investigated the use of LLMs together with REST APIs, it has been primarily focused on how to generate test cases for a OpenAPI specifications [26, 6]. In contrast, our goal is to generate code for applications that express their interfaces via REST APIs. Surprisingly, we do not know of any other works that address this problem.
In this paper we propose a complexity metric for MSBA specs, including dependency difficulty. Along somewhat similar lines, the authors in [32] define $6$ runnable levels to classify the scope required for a function to run successfully; e.g., from stand-alone functions to those that require knowledge of code in other source files. They also propose a metric Acc@k to measure the percent of target functions that are referenced in at least one of the $k$ LLM-generated sample codes. A similar measure, Recall@k, is defined in [34]. Our complexity metric is focussed on the specification complexity and includes additional components, such as length of the specification.
7 Conclusion
The goal of this paper is to evaluate how well LLMs do in generating code for microservice-based applications. We defined a standard template for specifying such applications, a framework for testing them, and proposed a metric for scoring the complexity of a specification; that is, the difficulty of generating code that satisfies the specification. We ran experiments on 8 microservices from 2 applications. Our results indicate, unsurprisingly, that the greater the complexity of the specification, the more difficulty the LLM has in passing unit tests for the microservices. While stronger LLMs such as GPT-3o-mini do fairly well in generating code for specifications of medium difficulty, they do very poorly on complex specifications that have more intricate business logic, more dependencies, and require authentication and database integration. We analyzed the key problems LLMs face in generating code for these specifications thereby suggesting future research directions to improve the capabilities of LLMs to generate microservice-based applications. It is yet to be seen whether stronger LLMs will require additional techniques to generate code for MSBAs or will be able to do so on their own.
Appendix A Prompts
The following prompt is used to generate code for the given microservice.
Below is (I) a description of the microservices for a public libraryapplication and (II) code generation guidelines. Generate codefor the <microservice-name>  microservice following the guidelines.The code you generate should be complete and executable.I.  <MSBA spec>II. <code generation guidelines>
This prompt provides the MSBA description and the code generation guidelines. The MSBA descriptions for the public library and the restaurant nutrition MSBAs are too long to include here. They can be found online (see footnote 2). The code generation guidelines are given in Figure 3. The reflection prompt used to regenerate code when there are errors is presented in Figure 4.
Code Generation Guidelines- Each microservice should be written in Python.- The Python packages that need to be installed   must be listed in a comment  at the beginning of the code.- The code must be executable. Besides comments in the code, no other text  in the response.- Each microservice should store its persistent data in a Mongo database.  Assume that Mongo is available at ’mongodb://mongo:27017/’.- If a microservice requires data that another microservice is responsible for,  it invokes a GET request on that microservice to retrieve  that data when it  needs it. It does not  replicate and store the data locally.- Assume that each microservice will run in its own Docker container.- For all services, each REST request should return a status code in addition  to the JSON payload if there is one.  For successful requests, the appropriate  2XX code is returned:  201 (successful POST), 204 (successful DELETE) and  200 (success on other requests).  For unsuccessful requests, the appropriate 4XX or 5XX code is returned: 400  (Bad request), 404 (Not Found), 405 (Method Not Allowed), 415 (Unsupported  media type), and 500 (Internal Server Error).
Figure 3: Code Generation Guidelines
Examine the error messages given above. For at most 5 errors, provide anexplanation of (a) where the faulty code is located, (b) what the erroris, and (c) how to fix it. Your explanation  should be precise so that aprogrammer can follow your instructions and correct the errors. Do notconsider Warnings or the fact that the container is not running on aproduction server. If there is not enough information in the errormessages, then consider which tests failed and examine the code to determinewhat changes to the code are needed. Think  what may have caused the error,including preceding requests. For example, a GET request may return a 404status code if the previous POST request adding that item failed.
Figure 4: The reflection prompt
The prompt for regenerating a program after reflection is:
"When the Python Interpreter tried to execute this program, there wereerrors.  Here is the summary and instructions on how to fix these errors:{reflection}.   Re-generate the code and make sure to fix the errors.Check that you have fixed each error."
Appendix B LLM-as-a-Judge
The LLM-as-a-Judge prompt is:
Measure the complexity of implementing the following micro-servicesapplication specification on a scale 1-5 (1 is easy, 5 is hard):<spec>The ratings and comments on the complexity of the MSBAs given in this paper are provided in Figures 5 and 6. We used Claude Sonnet 4 to provide these ratings. We note that we separately asked the same LLM to score each individual service. It was consistent for the restaurant app where it rated Authentication and Ratings at $2$ (out of $5$), the profile service at $3$ and the dishes service at $4$. However for the library app it rated Logs at $1$, Cardholders at $3$, Books at $4$, and Borrows at $5$.
Looking at this public library microservices application specification, Iwould rate the complexity as **3 out of 5** (moderate complexity).Here’s my analysis:**Factors that keep complexity moderate (not higher):**- **Clear, well-defined requirements** - The specification is detailed and    unambiguous- **Standard CRUD operations** - Most services follow basic Create, Read,    Update, Delete patterns- **Common technology stack** - Python, MongoDB, Docker, REST APIs are     well-established- **Simple data models** - Straightforward JSON documents with basic field     types- **Limited business logic** - Most complexity is in basic calculations    (fines, return dates)**Factors that add complexity:**- **Inter-service communication** - Services need to call each other    (Cardholders → Borrows, Borrows → Logs)- **External API integration** - Google Books API integration for    publisher/date lookup- **Business rule enforcement** - Preventing borrowing when 2+ books are    overdue- **Date calculations** - Fine calculations based on overdue days, return    date computation- **Error handling** - Proper HTTP status codes and exception handling across services- **Distributed data consistency** - Ensuring data integrity across multiple services    and databases**Why not higher complexity:**- No complex algorithms or data processing- No authentication/authorization requirements- No real-time or high-performance requirements- No complex integrations beyond one external API- No sophisticated business workflows**Why not lower complexity:**- Multiple interconnected services require coordination- Business logic beyond simple CRUD operations- External dependencies and error scenarios to handleThis falls squarely in the "moderate" category - more complex than a simple single-service CRUD app, but less complex than enterprise-scale distributed systems withcomplex business logic, security requirements, or performance constraints.
Figure 5: Claude Sonnet 4 comments on the complexity of the library MSBA.
Looking at this Restaurant-Nutrition microservices application specification, I wouldrate the implementation complexity as **4 out of 5 (Hard)**.Here’s my analysis of the complexity factors:**High Complexity Elements (Contributing to the 4 rating):**1. **Multi-service Architecture**: Four interconnected microservices requiring proper    service-to-service communication and coordination2. **Authentication & Authorization**: Token-based auth system with validation across    services, plus privileged user handling3. **External API Integration**: Integration with Ninja API for nutritional data,   including error handling for external service failures4. **Complex Business Logic**:   - Nutritional calculations (summing ingredient values)   - Diet matching algorithms (comparing dish nutrition to user profiles)   - Rating aggregation and average calculations5. **Database Operations**: Full CRUD operations across multiple MongoDB collections   with proper data modeling6. **Error Handling**: Comprehensive HTTP status code handling (401, 400, 404, 405,  415, 500) across all services7. **Query Processing**: Complex filtering logic for dishes (nutritional constraints,  ingredient matching) and profiles**Moderate Complexity Elements:**- Docker containerization requirements- REST API design and implementation- Data validation and formatting (2 decimal places for floats)**Why not a 5 (Extremely Hard):**- Well-defined specifications with clear requirements- Standard technologies (Python, MongoDB, REST APIs)- No real-time processing or complex distributed systems patterns- No advanced security beyond basic token authThis is a substantial project that would require solid understanding of microservicesarchitecture, database design, API integration, and comprehensive error handling,making it challenging but achievable for developers with intermediate to advanced skills.
Figure 6: Claude Sonnet 4 comments on the complexity of the restaurant MSBA.
Appendix C 1-shot prompt
Figure 7 presents the example used in our 1-shot prompt. For brevity, we only present the code for the POST request.
Here is example code for a REST service that supports POST, GET, DELETEand PUT requests on "toys" using Python, Flask and MongoDB. Each toy is aJSON object containing the fields name (of type string), descr (of type string),age (of type integer), price (of type float), features (of type array, whereeach element of the array is a string), and id (unique string id).from flask import Flask, jsonify, request, make_responseimport pymongofrom bson import ObjectIdimport osapp = Flask(__name__)client = pymongo.MongoClient(‘mongodb://mongo:27017/’)db = client[‘toys_inventory’]toysColl = db[‘toys’]@app.route(‘/toys’, methods=[‘POST’])def addToy():    try:        content_type = request.headers.get(‘Content-Type’)        if content_type != ‘application/json’:            return jsonify({"error": "Expected application/json"}), 415        data = request.get_json()        required_fields = [‘name’, ‘age’, ‘price’]        if not all(field in data for field in required_fields):            return jsonify({"error": "Malformed data"}), 400        if ‘features’ not in data:            features = []        else:            features = data[‘features’]        if ‘descr’ not in data:            descr = "Not Available"        else:            descr = data[‘descr’]        toy = {            ‘name’:data[‘name’],            ‘descr’:descr,            ‘age’:data[‘age’],            ‘price’:data[‘price’],            ‘features’:features            }        result = toysColl.insert_one(toy)        id = str(result.inserted_id)        response_data = {"id":id}        return jsonify(response_data),201    except Exception as e:        return jsonify({"server error":str(e)}),500
Figure 7: Example used in the 1-shot prompt
Appendix D The effects of temperature
To see the effects of temperature, we analyze GPT-o4-mini on the 0-shot prompt for the library MSBA. Table 8 presents the results. The meaning of the columns are the same as in other tables of Section 5. Each row gives the results for one temperature for one MS, and the row All is the average across all temperatures for that MS [^6] Note that the averages given for the All column cannot be obtained by simply adding up the the percent passed for the $3$ temperatures and then dividing by $3$. This is because to compute the average, we only consider those experiments that actually generated testable code. Hence to validate the results of the All row you must take a weighted sum of the $3$ temperatures where the weights are determined by the number of testable programs that were generated..
In this table, for V0 code, temperature $0.3$ outperforms the other temperatures except for one MS (Cardholder), where temperature 0 performs best. The difference between the top and second best performing temperatures are always within $10\%$, and for $2$ of the MSes (Borrows and Logs), all the temperatures are within $10\%$ of the top performing temperature. For V1 code temperature $0.5$ performs best across all MSes.
Name temp # V0 # V1 # V0 % V1
testable testable passed passed
Cardholder 0.0 5/5 5 77.1 80.0
Cardholder 0.3 4/5 3 48.2 52.4
Cardholder 0.5 4/5 4 69.6 82.1
Cardholder All 13/15 12 65.9 73.8
Books 0.0 1/5 1 35.7 35.7
Books 0.3 4/5 2 73.2 60.7
Books 0.5 4/5 3 71.4 73.8
Books All 9/15 6 68.2 63.1
Borrows 0.0 5/5 5 57.1 58.6
Borrows 0.3 5/5 5 61.4 68.6
Borrows 0.5 5/5 5 55.7 71.4
Borrows All 15/15 16 58.1 66.2
Logs 0.0 5/5 5 85.2 85.2
Logs 0.3 5/5 1 94.2 85.7
Logs 0.5 5/5 4 88.6 85.7
Logs All 15/15 10 89.4 85.5
Table 8: Effects of temperature: Library MSBA, GPT-4o-mini, 0-shot